show that u e^-2xy sin x2 y2 is harmonic

1
Expert's answer
2022-01-10T12:24:22-0500
"u=e^{-2xy}\sin(x^2-y^2)"

"\dfrac{\partial u}{\partial x}=-2ye^{-2xy}\sin(x^2-y^2)+2xe^{-2xy}\cos(x^2-y^2)"

"\dfrac{\partial^2 u}{\partial x^2}=4y^2e^{-2xy}\sin(x^2-y^2)-4xye^{-2xy}\cos(x^2-y^2)"

"+2e^{-2xy}\cos(x^2-y^2)-4xye^{-2xy}\cos(x^2-y^2)"

"-4x^2e^{-2xy}\sin(x^2-y^2)"

"\dfrac{\partial u}{\partial y}=-2xe^{-2xy}\sin(x^2-y^2)-2ye^{-2xy}\cos(x^2-y^2)"

"\dfrac{\partial^2 u}{\partial y^2}=4x^2e^{-2xy}\sin(x^2-y^2)+4xye^{-2xy}\cos(x^2-y^2)"

"-2e^{-2xy}\cos(x^2-y^2)+4xye^{-2xy}\cos(x^2-y^2)"

"-4y^2e^{-2xy}\sin(x^2-y^2)"


"\dfrac{\partial^2 u}{\partial x^2}+\dfrac{\partial^2 u}{\partial y^2}=4y^2e^{-2xy}\sin(x^2-y^2)"

"-8xye^{-2xy}\cos(x^2-y^2)+2e^{-2xy}\cos(x^2-y^2)"

"-4x^2e^{-2xy}\sin(x^2-y^2)+4x^2e^{-2xy}\sin(x^2-y^2)"

"+8xye^{-2xy}\cos(x^2-y^2)-2e^{-2xy}\cos(x^2-y^2)"

"-4y^2e^{-2xy}\sin(x^2-y^2)=0"

Hence


"\nabla^2u=\dfrac{\partial^2 u}{\partial x^2}+\dfrac{\partial^2 u}{\partial y^2}=0, x, y\in \R"

Therefore the function "u=e^{-2xy}\sin(x^2-y^2)" is harmonic.